### v1

# Coxeter embeddings are injective

Abstract.
We show that certain embeddings of Coxeter groups within other Coxeter groups are injective.
1. Coxeter partitions and Coxeter embeddings
Throughout, a Coxeter system will be denoted as $(W,S)$, where $W$ is a Coxeter group and $S$ is the corresponding set of Coxeter generators. For $J\subseteq S$, the corresponding parabolic subgroup of $W$ will be denoted by $W_{J}$. Given words $u$ and $w$, and $m\geq 2$, the notation $\langle u,w\rangle^{m}$ denotes the $m$-alternating expression:
$\langle u,w\rangle^{m}:=\underbrace{uwu\ldots}_{m\text{ times}}.$
A subword is always assumed to be contiguous.
A surjective function $\pi\colon\hat{S}\to S$ will partition the set $\hat{S}$ into a disjoint union of preimages $U(s):=\pi^{-1}(s)$.
Definition 1.1.
Let $(\hat{W},\hat{S})$ be a Coxeter system, where $\hat{S}$ is finite, and let $S$ be a finite set. A Coxeter partition is a surjective function $\pi\colon\hat{S}\to S$ such that
  1. (1)
    For each $s\in S$, the finite parabolic subgroup $\hat{W}_{U(s)}$ associated to $U(s)\subseteq\hat{S}$ is a finite Coxeter group. Let $w_{0}^{s}$ denote the longest element of $\hat{W}_{U(s)}$.
  2. (2)
    For each pair of elements $s\neq t\in S$, we denote the order of the element $w_{0}^{s}w_{0}^{t}\in\hat{W}$ by $m_{st}$ (possibly infinite). For some (and hence every) reduced expression of $w_{0}^{s}$ and $w_{0}^{t}$, we require the following:
    1. (a)
      If $m_{st}<\infty$, then the parabolic subgroup associated to $U(s)\sqcup U(t)$ is finite, and $\langle w_{0}^{s},w_{0}^{t}\rangle^{m_{st}}$ is a reduced expression of its longest element.
    2. (b)
      If $m_{st}=\infty$, then for all $k\geq 2$, we require that $\langle w_{0}^{s},w_{0}^{t}\rangle^{k}$ is a reduced expression.
Given a Coxeter partition of $(\hat{W},\hat{S})$, let $(W,S)$ be the Coxeter system with simple reflections $S$, where $st$ has order $m_{st}$ for each $s\neq t\in S$.
We note here that condition 2(a) implies that $\langle w_{0}^{s},w_{0}^{t}\rangle^{k}$ is reduced for each $2\leq k\leq m_{st}$, being a subword of the reduced expression $\langle w_{0}^{s},w_{0}^{t}\rangle^{m_{st}}$.
Definition 1.2.
The Coxeter embedding (associated to the Coxeter partition $\pi$) is the homomorphism $\phi\colon W\to\hat{W}$ defined on generators by $s\mapsto\phi(s):=w_{0}^{s}$.
Proposition 1.3.
The Coxeter embedding is a well-defined homomorphism.
Proof.
For $s\neq t\in S$, the element $\phi(s)\phi(t)=w_{0}^{s}w_{0}^{t}$ has order $m_{st}$ by definition. Any longest element is an involution, so $\phi(s)^{2}=\operatorname{id}$ for any $s\in S$. ∎
Our main theorem is as follows:
Theorem 1.4.
Coxeter embeddings are injective, and send reduced expressions to reduced expressions. Moreover, for $w\in W$ and $\hat{s}\in\hat{S}$, letting $s=\pi(\hat{s})$, we have $\phi(w)\hat{s}<\phi(w)$ if and only if $ws<w$.
The rest of this section is devoted to proving this theorem.
We will use some standard notions and results for Coxeter groups, which can all be found (or deduced) from [BB05, §3 and §4]. From here on, the partial order $\leq$ (and $<$) on any Coxeter system refers to the weak right Bruhat order. We use the following notation. For elements $x,y,z$ in a Coxeter group, we write $x=y.z$ whenever $x=yz$ and $\ell(x)=\ell(y)+\ell(z)$; this is called a reduced composition. We write $\operatorname{RD}(w)$ for the right descent set of $w$, i.e. the set of simple reflections $s\in S$ such that $ws<w$. Recall that $\operatorname{RD}(w)$ always generates a finite Coxeter group. Moreover, if $x=y.z$ then $\operatorname{RD}(z)\subset\operatorname{RD}(x)$.
If $(W,S)$ is a Coxeter system and $J\subseteq S$, write $W^{J}$ for the set of minimal length representatives for cosets in $W/W_{J}$. If $y\in W^{J}$ and $z\in W_{J}$ then $y.z$ is reduced. Any $x\in W$ has a unique decomposition $x=y.z$ with $y\in W^{J}$ and $z\in W_{J}$.
Lemma 1.5.
Let $(W,S)$ be any Coxeter system and $J\subseteq S$. Suppose we have $a\in J$, $B\in W_{J}$ and $C\in W$ such that $C.B$ is reduced but $CBa$ is not. Then either $a\in\operatorname{RD}(B)$ or $\operatorname{RD}(C)\cap J\neq\emptyset$.
Proof.
Suppose $\operatorname{RD}(C)\cap J=\emptyset$. This implies $C\in W^{J}$. Since $Ba\in W_{J}$, $C.(Ba)$ is reduced. If $Ba$ were reduced, then $CBa$ would be reduced. Thus $Ba$ is not reduced, implying $a\in\operatorname{RD}(B)$. ∎
Proof of Theorem 1.4.
We show that a Coxeter embedding $\phi$ sends reduced expressions of $w\in W$ to reduced expressions in $\hat{W}$, by induction on the length of $w$. For $\ell(w)\leq 1$, the statement follows from the definition of Coxeter embeddings. From now on let $w\in W$ with $\ell(w)\geq 2$, and assume via induction that the statement holds for all reduced expressions of elements in $W$ with length $<\ell(w)$.
Pick a reduced expression for $w$, and let $s$ and $t$ be the last two letters of this reduced expression. Let $J=\{s,t\}\subset S$, and consider the unique decomposition $w=x.y$ with $x\in W^{J}$ and $y\in W_{J}$. Note that $\ell(y)\geq 2$. As such, $\ell(x)\leq\ell(w)-2$ and so $\phi(x)$ is reduced by the inductive hypothesis. Moreover, as an element of the dihedral group $W_{J}$, $y$ is an alternating product of $s$ and $t$, so Condition (2) of Coxeter partitions guarantees that $\phi(y)$ is reduced. The statement is proven if $\phi(x).\phi(y)$ is reduced.
Assume to the contrary that $\phi(x)\phi(y)$ is not reduced. We decompose the reduced expression $\phi(y)\in\hat{W}_{U(s)\sqcup U(t)}$ into $B.a.A$ with $a\in U(s)\sqcup U(t)\subseteq\hat{S}$ so that $\phi(x).B$ is reduced but $\phi(x)Ba$ is not. Note that $B.a$ is reduced by construction, so $a\not\in\operatorname{RD}(B)$. By Lemma 1.5, there exists some $b\in\operatorname{RD}(\phi(x))\cap(U(s)\sqcup U(t))\neq\emptyset$. Suppose $b\in U(s)$. Then $b\leq w_{0}^{s}=\phi(s)$ and so $\phi(xs)=\phi(x)\phi(s)$ is not reduced. But $xs=x.s$ is reduced (since $x\in W^{s,t}$) and $\ell(xs)=\ell(x)+1<\ell(w)$. By the induction hypothesis $\phi(x)\phi(s)$ must be reduced, which is a contradiction. The case $b\in U(t)$ is treated in exactly the same fashion.
Since reduced expressions are sent to reduced expressions, $\phi$ has trivial kernel.
We now prove the final statement of the theorem. An equivalent statement is that either $s\in\operatorname{RD}(w)$ and $U(s)\subset\operatorname{RD}(\phi(w))$, or $s\notin\operatorname{RD}(w)$ and $U(s)\cap\operatorname{RD}(\phi(w))=\emptyset$. Suppose $ws>w$. Then $\phi(w)\phi(s)$ is reduced so $\operatorname{RD}(\phi(w))\cap U(s)=\emptyset$. Conversely, suppose that $ws<w$. Then $w$ has a reduced expression ending in $s$, so $\phi(w)$ has a reduced expression ending in $\phi(s)$. Thus $U(s)\subset\operatorname{RD}(\phi(w))$. ∎
Remark 1.6.
The proof of Theorem 1.4 is modelled on [Cri99, Lemma 2.2]. However, we note that Lemma 2.2, as stated in loc. cit., is actually false. Namely, an LCM-homomorphism $\phi:A^{+}_{\Gamma}\rightarrow A^{+}_{\hat{\Gamma}}$ does not guarantee that $\phi$ sends square-free elements to square-free elements (which allows one to deduce that the induced map on the corresponding Coxeter groups sends reduced expressions to reduced expressions). Consider the Artin monoids associated to the following Coxeter graphs:
{tikzpicture}
The homomorphism defined by $\phi(s)=s_{1}s_{2}$ and $\phi(t)=t_{1}t_{2}$ is an LCM-homomorphism, but it does not send square-free elements to square-free elements, since
$\phi(sts)=s_{1}s_{2}t_{1}t_{2}s_{1}s_{2}=s_{2}t_{1}s_{2}t_{2}s_{1}s_{1}$
is not square-free. The issue is that when $st\in A^{+}_{\Gamma}$ has no common multiple, an LCM-homomorphism imposes no condition on the images of $\langle s,t\rangle^{m}$. Condition 2(b) in our definition of Coxeter partitions was added to handle exactly this situation.
2. Examples of Coxeter embeddings
We use the convention that an irreducible infinite type Coxeter system $(W,S)$ has Coxeter number $\infty$.
We present here some examples of Coxeter partitions of interest, which include (and generalize) some examples that were considered by others.
Definition 2.1.
Let $(\hat{W},\hat{S})$ and $(W,S)$ be two Coxeter systems, with $\hat{S}$ and $S$ finite sets. A Lusztig’s partition is a surjective function $\pi\colon\hat{S}\to S$ such that
  1. (1)
    For each $s\in S$, the elements of $U(s)$ all commute with each other.
  2. (2)
    For each pair $s\neq t\in S$, the parabolic subgroup $\hat{W}_{U(s)\sqcup U(t)}$ generated by $U(s)\sqcup U(t)$ is a product of Coxeter systems which all have the same Coxeter number $m_{st}$ (whether finite or infinite).
Said another way, a Lusztig’s partition is a coloring of the vertices $\hat{S}$ in the Coxeter graph of $\hat{W}$, such that no color is adjacent to itself, and such that any two colors form a disjoint union of Coxeter graphs with the same Coxeter number.
It is well-known when $\hat{W}$ is finite that the morphism $\phi:W\to\hat{W}$ associated to a Lusztig’s partition defined by
$\phi(s)=\prod_{\hat{s}\in U(s)}\hat{s}$
sends reduced expressions to reduced expressions (e.g. via Lemma 2.2 below). We prove that this remains true even when $\hat{W}$ is infinite, by showing that Lusztig’s partitions are a specific type of Coxeter partitions. For this we will require two results about reduced expressions for powers of Coxeter elements. The first lemma is classical, and the second is a result of Speyer.
Lemma 2.2 ([Bou02, Chapter V §6 Exercise 2]).
Let $(W,S)$ be an irreducible finite Coxeter system with Coxeter number $h$. Let $S=S^{\prime}\sqcup S^{\prime\prime}$ be a partition of $S$ so that in each subset all elements commute. Denote $x:=\prod_{s^{\prime}\in S^{\prime}}s^{\prime}$ and $y:=\prod_{s^{\prime\prime}\in S^{\prime\prime}}s^{\prime\prime}$. Then the longest element $w_{0}$ is given by
$w_{0}=\langle y,x\rangle^{h}=\langle x,y\rangle^{h}$
and both expressions are reduced. In particular, the largest power of the Coxeter element $c:=x.y$ that is reduced is $c^{\lfloor h/2\rfloor}$.
Lemma 2.3 ([Spe09, Theorem 1]).
Let $(W,S)$ be an irreducible infinite Coxeter system. For any choice of Coxeter element $c\in W$, $c^{k}$ is a reduced expression for all $k$.
Proposition 2.4.
Lusztig’s partitions are Coxeter partitions. Conversely, a Coxeter partition where the elements of $U(s)$ all commute with each other (for each $s\in S$) is a Lusztig’s partition.
Proof.
Given a surjection $\pi:\hat{S}\rightarrow S$ satisfying condition (1) of a Lusztig’s partition, it is immediate that condition (1) of a Coxeter partition is satisfied. It suffices to prove that condition (2) of a Lusztig’s partition is equivalent to condition (2) of a Coxeter partition, in this case. Condition (2) is a “rank two condition,” namely it is a condition on each pair of distinct vertices $s\neq t\in S$. As such, it suffices to prove the case where $S=\{s,t\}$ is a two-element set. This will be the case considered in the rest of this proof.
First suppose that $(\hat{W},\hat{S})$ corresponds to a connected Coxeter graph. Then $\hat{S}$ has a unique bipartite coloring (up to swapping the colors). Condition (2) of a Lusztig’s partition is now vacuous, so we must prove that condition (2) of a Coxeter partition holds. If $\hat{W}$ is finite with Coxeter number $m_{st}$, that $\langle w_{0}^{s},w_{0}^{t}\rangle^{m_{st}}$ is a reduced expression for the longest element of $\hat{W}$ follows from Lemma 2.2. If $\hat{W}$ is instead infinite, consider the Coxeter element $\hat{c}$ of $\hat{W}$ given by
$\hat{c}:=w_{0}^{s}w_{0}^{t}=\left(\prod_{\hat{s}\in U(s)}\hat{s}\right)\left(\prod_{\hat{t}\in U(t)}\hat{t}\right).$
By Lemma 2.3, any power of $\hat{c}$ is reduced [^1] In fact, we only need a weaker version of Lemma 2.3, which was proven earlier in [FZ07, Corollary 9.6].. But for all $k\geq 2$, $\langle w_{0}^{s},w_{0}^{t}\rangle^{k}$ is always a subword of some power of $\hat{c}$, so $\langle w_{0}^{s},w_{0}^{t}\rangle^{k}$ must itself be reduced.
Now consider the general case: $\hat{S}$ is a disjoint union of connected Coxeter graphs $\hat{S}_{i}$. Then $\hat{W}$ is a product $\prod_{i}\hat{W}_{i}$. For each $i$, we use the shorthand
$w_{0}^{s,i}:=\prod_{\hat{s}\in U(s)\cap\hat{S}_{i}}\hat{s},$
so that $w_{0}^{s}$ is the product of all $w_{0}^{s,i}$ (in any order). For all $k\geq 2$, the following equality (in $\hat{W}$) can be obtained by only applying commutativity relations:
$\langle w_{0}^{s},w_{0}^{t}\rangle^{k}$ $=\left\langle\prod_{i}w_{0}^{s,i},\prod_{i}w_{0}^{t,i}\right\rangle^{k}$
$=\prod_{i}\langle w_{0}^{s,i},w_{0}^{t,i}\rangle^{k}.$
If each $\hat{W}_{i},$ has the same Coxeter number $m_{st}$ (finite or infinite), then by the connected case above, $\langle w_{0}^{s,i},w_{0}^{t,i}\rangle^{k}$ is reduced for $2\leq m\leq m_{st}$. As a (direct) product of reduced expressions, $\langle w_{0}^{s},w_{0}^{t}\rangle^{k}$ is also reduced, which is condition (2) of a Coxeter partition.
Conversely, let $m_{st}$ be the order of $w_{0}^{s}w_{0}^{t}$ (finite or infinite). If some $\hat{W_{i}}$ has Coxeter number $h_{i}$ strictly less than $m_{st}$, then for any $h_{i}<k\leq m_{st}$ we have that $\langle w_{0}^{s,i},w_{0}^{t,i}\rangle^{k}$ is not a reduced expression. Consequently, neither is $\langle w_{0}^{s},w_{0}^{t}\rangle^{k}$. So if condition (2) of a Coxeter partition holds, then $h_{i}\geq m_{st}$ whenever $m_{st}$ is finite, or $h_{i}=\infty$ whenever $m_{st}$ is infinite. Since $h_{i}$ divides $m_{st}$ whenever both are finite (in order for $(w_{0}^{s}w_{0}^{t})^{m_{st}}$ to be the identity), we deduce that $h_{i}=m_{st}$. This proves condition (2) of a Lusztig’s partition. ∎
Condition (2) of Lusztig’s partitions is a “same Coxeter number” property for connected components of subgraphs generated by $U(s)\sqcup U(t)$. We now show that the “same Coxeter number” property extends to connected components of the whole Coxeter graph222 By restricting the Coxeter partition from $\hat{S}$ to a suitable subset, we also obtain the result for $U(s)\sqcup U(t)\sqcup U(u)$ and other preimages. of $(\hat{W},\hat{S})$. Note that this is not true for Coxeter partitions in general; see Example 2.12.
Proposition 2.5.
Let $\pi:\hat{S}\rightarrow S$ be a Lusztig’s partition and let $(W,S)$ be an irreducible Coxeter system with Coxeter number $h$. Then the irreducible components of $(\hat{W},\hat{S})$ also have the same Coxeter number $h$.
Proof.
Throughout, we will implicitly use the fact that a Lusztig’s partition is a Coxeter partition, shown in Proposition 2.4.
Since $\pi$ is a Lusztig’s partition, the associated Coxeter embedding $\phi:W\rightarrow\hat{W}$ sends any Coxeter element $c\in W$ to a (mutually commuting) product $\prod_{i}\hat{c}_{i}$ of Coxeter elements $\hat{c}_{i}$ for each irreducible component $(\hat{W}_{i},\hat{S}_{i})$ of $(\hat{W},\hat{S})$. By Theorem 1.4, $\phi$ sends reduced expressions to reduced expressions, hence if $c^{k}$ is reduced, then so is $\prod_{i}(\hat{c}_{i})^{k}$. Since the $\hat{c}_{i}$’s mutually commute, the final statement is equivalent to each $(\hat{c}_{i})^{k}$ being reduced.
Suppose $h=\infty$. By Lemma 2.3, $c^{k}$ is reduced for all $k>0$, and thus so is $(\hat{c}_{i})^{k}$. This shows that each irreducible component $(\hat{W}_{i},\hat{S}_{i})$ is infinite type, as desired.
Now let us assume $h<\infty$, and prove that $h_{i}=h$, where $h_{i}$ is the Coxeter number of the component $(\hat{W}_{i},\hat{S}_{i})$. Take the Coxeter element $c=x.y\in W$ associated to some (two-coloring) partition of $S$ as in Lemma 2.2, so that $c^{\lfloor h/2\rfloor}$ is a reduced expression. By Theorem 1.4, we get that
$\phi(c^{\lfloor h/2\rfloor})=\prod_{i}(\hat{c}_{i})^{\lfloor h/2\rfloor}$
is a reduced expression, and thus each $(\hat{c}_{i})^{\lfloor h/2\rfloor}$ is a reduced expression. In particular, the order $h_{i}$ of $\hat{c}_{i}$ satisfies $h_{i}>\lfloor h/2\rfloor$. Since $(\hat{c}_{i})^{h}=\operatorname{id}$, $h_{i}$ divides $h$. A simple numerical argument shows that $h_{i}=h$ as required. ∎
Coxeter embeddings associated to Lusztig’s partitions (when each $m_{st}$ is finite) were studied by Lusztig in [Lus83, §3]. These include examples coming from folding by graph symmetries, such as the first example below. Lusztig’s famous inclusion of Coxeter groups of type $H_{4}$ into type $E_{8}$ is a Lusztig’s partition that does not come from graph symmetries. Our setting of Lusztig’s partitions includes examples where $m_{st}$ is infinite (see Example 2.8).
We present below some examples of Lusztig’s partitions, followed by some examples of Coxeter partitions which are not Lusztig’s partitions. In all examples to follow, the elements in $\hat{S}$ are given by alphabets with subscripts and elements in $S$ are given by alphabets without subscripts. The partition map $\pi:\hat{S}\rightarrow S$ is defined by forgetting the subscript.
Example 2.6.
Consider the folding of $D_{4}$ onto $G_{2}$.
{tikzpicture}
This is an example of folding from a graph symmetry, because the map $\pi$ records the orbits under a group action on the Coxeter graph.
Example 2.7.
Here are two different Lusztig’s partitions, giving embeddings of $G_{2}$ which do not come from graph symmetries.
{tikzpicture}
Example 2.8.
Let $\hat{\Gamma}$ be a bipartite Coxeter graph, all of whose connected components have the same Coxeter number $h$ (possibly infinite). Any two coloring of $\hat{\Gamma}$ defines a Lusztig’s partition onto the dihedral group $I_{2}(h)$. Below is an example for $h=\infty$.
{tikzpicture}
Example 2.9.
The following is an example of a Lusztig’s partition similar to the Lusztig’s partition from $E_{8}$ onto $H_{4}$, except with type $B$ and type $D$ tails.
{tikzpicture}
Example 2.10.
This example takes an embedding of $I_{2}(12)$ into $E_{6}\times I_{2}(12)$, and adds some fun.
{tikzpicture}
We conclude with some examples of Coxeter partitions that are not Lusztig’s partitions.
Example 2.11.
Consider the following partition of $A_{4}$ onto $B_{2}$:
{tikzpicture}
This is not a Lusztig’s partition since $U(t)=\{t_{1},t_{2}\}$ and $t_{1}t_{2}\neq t_{2}t_{1}$. (Note that $A_{4}$ and $B_{2}$ have different Coxeter numbers; cf. Proposition 2.5.) Nonetheless, $U(s)$ and $U(t)$ both generate a finite parabolic subgroup, so condition (1) of a Coxeter partition holds. The corresponding longest elements are $w_{0}^{s}=s_{1}s_{2}=s_{2}s_{1}$ and $w_{0}^{t}=t_{1}t_{2}t_{1}=t_{2}t_{1}t_{2}$ respectively. A straightforward calculation shows that
$\langle w_{0}^{s},w_{0}^{t}\rangle^{4}=s_{2}s_{1}t_{2}t_{1}t_{2}s_{2}s_{1}t_{2}t_{1}t_{2}$
is a reduced expression for the longest element of $A_{4}$. The order of $w_{0}^{s}w_{0}^{t}$ is indeed $4$, since its square is the longest element and has order $2$. This shows that condition (2) is satisfied. Similarly, we have a Coxeter partition (that is not a Lusztig’s partition) of affine type $D_{5}$ onto affine type $C_{2}$:
{tikzpicture}
The reader is encouraged to construct the affine-type examples given in [Cri99, Table 2].
Example 2.12.
Note that Coxeter partitions (but not Lusztig’s partitions; see Proposition 2.5) allow mixed Coxeter numbers, as the following example from $A_{3}\sqcup A_{4}$ to $B_{2}$ shows:
{tikzpicture}